OpenStudy (anonymous):

Solve operation with complex numbers: [ (1 + i) / i ] + [ -3 / (4-i) ]. Is the answer [ (-1 + i) / -1 ] + [ (-12 - 3i) / 17 ]?

OpenStudy (anonymous):

common denominator: \[[(1+i)(4-i) -3*i]/(i*(4-i))=(4+4i-i+1-3i)/(4i+1)=5/(4i+1)\] \[=5/(4i+1)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I know this is really stupid of me, but I can't see how it goes along with the original equation. Is there anyway you can show me that, all together?